OpenStudy (kmcunningham):

If y = cos^2 (2x^2) , find y'.

OpenStudy (anonymous):

\[y=\cos ^2\left( 2x^2 \right),y \prime=2\cos 2x^2*(-\sin 2x^2)*4x=-4x \sin 4x^2\]

OpenStudy (kmcunningham):

So it's 2 cos 2x?

OpenStudy (kmcunningham):

*2x^2? Sorry I'm very confused

OpenStudy (kmcunningham):

Oh wait I see the rest of it Nevermind thank you

OpenStudy (anonymous):

\[\frac{ d }{ dx }x^n=nx ^{n-1}\] \[\frac{ d }{ dx }\cos x=-\sin x\]

OpenStudy (anonymous):

instead of x here \[2x^2\]